If volume of mixture co and co2 increases by 20% on passing over hot coke, the percentage of co2 in the oeiginal mixture
1
Expert's answer
2018-08-06T07:44:09-0400
Answer:
If we have 100 ml CO and CO2, after reaction 120 ml CO (because CO2 + C = 2CO, 100*0.2=120)
nCO2=X ml
nCO=Y ml
CO2+C=2CO
2X-X=20 (because 2-1)
X=20
nCO2=20 ml
nCO=100-20=80 ml
So the percentage of co2 in the oeiginal mixture 80%, because (80*100)/100=80 %